Ryan has a bag of marbles with 4 blue marbles, 3 white marbles, and 6 red marbles.
Find the following probabilities of Ryan drawing the given marbles from the bag if the first marble(s) is (are) not returned to the bag after they are drawn.
(Answer as a reduced fraction)
a A Blue, then a Red = ________
b A Red, then a White = _______
c A Blue, then a Blue, then a Blue = _______

Answer by math_helper(2461)

Total number of marbles: 4+3+6 = 13

a) There are 4 outcomes (of 13 possible) for choosing blue: 4/13 for choosing blue.
That leaves 12 marbles in the bag, of which 6 are red, so 6/12 for choosing red on the 2nd choice.


P(blue then red) = (4/13)*(6/12) = 2/13

Parts (b) and (c) are solved in a similar manner.